### v1

# Integrating Neural Operators with Diffusion Models Improves Spectral Representation in Turbulence Modeling

Abstract
We integrate neural operators with diffusion models to address the spectral limitations of neural operators in surrogate modeling of turbulent flows. While neural operators offer computational efficiency, they exhibit deficiencies in capturing high-frequency flow dynamics, resulting in overly smooth approximations. To overcome this, we condition diffusion models on neural operators to enhance the resolution of turbulent structures. Our approach is validated for different neural operators on diverse datasets, including a high Reynolds number jet flow simulation and experimental Schlieren velocimetry. The proposed method significantly improves the alignment of predicted energy spectra with true distributions compared to neural operators alone. Additionally, proper orthogonal decomposition analysis demonstrates enhanced spectral fidelity in space-time. This work establishes a new paradigm for combining generative models with neural operators to advance surrogate modeling of turbulent systems, and it can be used in other scientific applications that involve microstructure and high-frequency content. See our project page: vivekoommen.github.io/NO_DM
Keywords Neural Operators  $\cdot$ Diffusion Models  $\cdot$ Turbulence Modeling
1 Introduction
Understanding turbulence is the last frontier in classical Physics. The complexity arises due to the high Reynolds numbers encountered in atmospheric and engineering turbulence, which induce an energy cascade producing spatiotemporal scales spanning multiple orders of magnitude, rendering them difficult to resolve fully. Also, extreme sensitivity to important other parameters, such as body roughness and external flow perturbations have a compound effect. In 1972, Howard Emmons, a distinguished researcher from Harvard, famously asserted that turbulence would never be computable. However, with the advent of supercomputers and developing dealiased spectral methods, a breakthrough occurred just a couple of years later. Steven Orszag of MIT, utilizing the NCAR’s first supercomputer (CDC7600), conducted the first direct numerical simulation (DNS) of homogeneous turbulence, albeit at lower Reynolds numbers (Re).
Progress has been made since then on simulating accurately several fundamental turbulent flows [1, 2, 3] but even today with exascale supercomputers and a speed-up of over $10^{10}$ over the CDC7600, DNS of turbulence is still limited to relatively low Reynolds numbers and simple-geometry flows. On the experimental side, methods such as particle image velocimetry (PIV) and particle tracking velocimetry (PTV) can provide useful information but cannot lead to quantitative understanding as DNS does, e.g. compute all critical statistical quantities of interest, and they may have difficulty resolving high wavenumbers and high frequencies.
It is clear, therefore, that a fundamentally different approach is needed to break the deadlock. What is new now that was not available before is the in-situ diagnostic tools, such as PIV and PTV, that can provide at least some partial information about the flow dynamics both in time and space. However, existing simulation methods are based on conventional numerical solvers, e.g. spectral methods, that require elaborate data assimilation techniques to integrate measurements. To this end, the recent emergence of physics-informed neural networks (PINNs) [4, 5, 6] enable seamless integration of multimodal data, and offer new opportunities for drastically advancing turbulence modeling. Indeed, PINNs have been very effective in discovering new physics even from sparse data by employing the so-called “gray-box" models. In this new paradigm, we postulate a mathematical model, e.g., the Navier-Stokes equations (NSE) but with an unknown correction term represented by a neural network (NN), whose weights and biases can be computed using the available data. Subsequently, the discovered NN can be expressed analytically using modern symbolic regression techniques, e.g. see [7, 8, 9, 10, 11], and new parameterization using fractional calculus to account for non-local interactions [12].
Another approach in assimilating data and provide faster solutions than DNS, which is prohibitively expensive at high Reynolds numbers, is through neural operators. In particular, classical DNS provides the solution for a given set of conditions. If there is a change in one of the given conditions, the solvers have to re-run which further adds to the computational cost. To circumvent this issue, neural operators were developed to work with a plurality of operating conditions and even different parameters, e.g. the Reynolds number $Re$ [13, 14, 15, 16, 17, 18]. In a neural operator the objective is to learn the mapping between two or more infinite-dimensional functional spaces. Although training a neural operator offline is expensive, the trained neural operator can infer solutions of unseen conditions almost instantaneously online, with negligible computational costs. Hence, neural operators can serve as efficient surrogates for different applications [19, 20, 21, 22, 23, 16, 24, 25].
Despite the promising results, inference of nonlinear dynamics and its temporal evolution using neural operators suffer from over-smoothing and loss of its high frequency content, a critical property of turbulent flows. This can also be attributed to the spectral bias of neural networks, which is the behavior of neural networks to learn only the low frequencies while remaining oblivious to higher frequency components. The spectral bias of neural operators/networks becomes increasingly significant in the surrogate modeling of turbulent systems whose energy spectrum decays rather slowly leading to non-trivial energy levels at higher wave numbers. In recent literature, the spectral bias has been addressed in [26] who used iterative refinements and [27] who blended neural operators with relaxation methods to alleviate the spectral bias of neural operators. Other ideas implemented for PINNs can be found in [28, 29].
In the current work, we investigate the use of generative AI models to overcome the spectral bias barrier. Surveying the recent advances in generative modeling demonstrates that generative adversarial networks (GANs) [30], normalizing flows [31] or diffusion models [32, 33] are capable of generating high-quality images with spectral properties that are desirable in turbulence modeling. Specifically, herein we aim to utilize the diffusion models to mitigate the spectral bias in neural operators. Diffusion models use gradual denoising of random noise to generate samples from the target distribution. Since diffusion models reverse the corruption process one step at a time, during the generative process they predict the next higher frequency based on the previously generated lower-frequency component. This is analogous to auto-regression in frequency space, and the generated images demonstrate better alignment across the frequency spectrum.
Recently, diffusion models have seen rapid development. Ho et al. [33] introduced the Denoising Diffusion Probabilistic Models (DDPM), which involves corruption of the signal with Gaussian noise followed by sequential denoising for retrieving signal at least from within the same distribution. Although DDPMs achieve high sample quality, the computational cost is high due to its requirement of a large number of sampling steps, typically 1000. Song et al. [34] demonstrated improved training and sampling by using neural networks to accurately estimate the score, which is a denoising-time-dependent gradient field of the perturbed data distribution, and then using numerical stochastic differential equation solvers such as Langevin MCMC (Markov Chain Monte Carlo) to generate samples. Karras et al. [35], performed a systematic investigation that helped elucidate the training and design of diffusion models by analyzing all types of diffusion models from within a common framework. The noise-weighted loss during training and improved sampling routines using preconditioned networks helped to bring down the number of sampling steps from $\approx 1000$ in DDPM to $\approx 30$. Furthermore, [36] demonstrated the efficiency of diffusion models on a variety of image-to-image tasks. For example, in the image colorization problem, a grayscale image serves as a prior for the diffusion model that generates the colorized version. Moreover, in the image inpainting challenge, the masked version of the image serves as an effective prior to the diffusion model that generates the full image. Similarly, [37, 38, 39] used statistically conditioned diffusion models for generating heterogeneous and statistically diverse microstructures. [40] builds stochastic and non-local closure models using conditional diffusion models and neural operators. The works in [41, 42] also present some of the recent applications of diffusion models on scientific data.
The ability of diffusion models to generate images with a slowly decaying energy spectrum, and the efficient utilization of priors motivated us to investigate the effectiveness of training a diffusion model conditioned on the neural operator’s output as their prior. First, we demonstrate that the neural operators suffer from spectral bias by illustrating predictions with over-smoothening. We mitigate the over-smoothening issue using the score-based diffusion model conditioned on the neural operator’s output. The improved spectral properties can be directly observed in our visualizations and the energy spectra plots, where we compare the ground truth and the neural operator’s predictions with the output of the conditioned diffusion models. Our approach is shown to work across a variety of neural operator architectures. We demonstrate the effectiveness of our framework on test cases with Reynolds numbers ranging from 2000 to $10^{6}$, 2D & 3D large-eddy simulations (LES), and raw experimental Schlieren velocimetry datasets. Furthermore, we perform a proper orthogonal decomposition (POD) analysis to provide a quantitative comparison of the energy decay and the POD modes of the predictions with the ground truth. Our ultimate objective is to accelerate DNS and LES by interweaving bursts of DNS/LES with the new combined method as well as to infer turbulent flow evolution at as DNS-level fidelity based on sparse experimental data.
2 Results
In the context of modeling physical systems, a non-linear mathematical operator $\bm{\mathcal{N}}$ that governs the evolution of a spatio-temporal field $\bm{u}(\bm{x},t)$, can be expressed as,
$\bm{u}(\bm{x},t+\Delta t)=\bm{\mathcal{N}}(\bm{u}(\bm{x},t))$ (1)
Unlike the conventional discretization-based solvers, neural operators ($\bm{\mathcal{G}}$) that learn to approximate the underlying true operator ($\bm{\mathcal{N}}\approx\bm{\mathcal{G}}$) from its solutions, have been shown to generalize well across the unseen initial conditions.
However, the neural operators often learn only the low wave numbers and ignore the higher wave numbers leading to an overall smoother approximation of the true solution. The issue is rooted in the spectral bias associated with the conventional training of the neural networks. In this section, we present the results of the experiments we performed to investigate whether diffusion models can help mitigate the issue of spectral bias in neural operators. The detailed methodology is explained in section 4 and is illustrated in Figure 1. All the neural operators and diffusion models used in this study have approximately 2 to 3 million trainable parameters.
Fig.  1: Methodology. First, we train a neural operator to learn mapping between the function spaces by minimizing, typically, the $L^{2}$ norm with respect to available ground truth. Next, we train a diffusion model conditioned on a neural operator’s output to approximate the ground truth data distribution using denoising score matching.
2.1 Kolmogorov Flow
The unsteady two-dimensional incompressible Navier-Stokes equation for a viscous, incompressible fluid with Kolmogorov forcing [43] in the vorticity form is given by:
$\begin{cases}\partial_{t}\omega+\bm{u}\cdot\nabla\omega=\nu\Delta\omega+f(x,y),&(x,y)\in(0,2\pi)^{2},t\in(0,t_{final}]\\ f(x,y)=\chi(\sin(2\pi(x+y))+\cos(2\pi(x+y))),&(x,y)\in(0,2\pi)^{2}\\ \nabla\cdot\bm{u}=0,&(x,y)\in(0,2\pi)^{2},t\in(0,t_{final}]\\ \omega(x,y,0)=\omega_{0},&(x,y)\in(0,2\pi)^{2}\end{cases}$ (2)
where $\chi\!=\!0.1$, $\omega$ is the vorticity, $\bm{u}$ is the velocity field vector, $\nu(=10^{-5})$ is the kinematic viscosity, and $\Delta$ is the two-dimensional Laplacian operator. We consider periodic boundary conditions. The source term $f$ is given by $f(x,y)=0.1(\sin(2\pi(x+y))+\cos(2\pi(x+y)))$, and the initial condition $\omega_{0}(x)$ is sampled from a Gaussian random field according to the distribution $\mathcal{N}(0,14^{1/2}(-\Delta+196I)^{-1.5})$. We used the publicly available pseudo-spectral solver [14] to generate a dataset with 1000 samples, and randomly split it to train, validation and test sets in the ratio 80:10:10. The dataset originally generated with a spatial resolution of $512\times 512$, was downsampled to $128\times 128$ to train the neural operators and diffusion models.
We train the neural operator ($\mathcal{G}$), to learn the mapping $\omega(x,y,t)_{|t\in[0,10]}\rightarrow\omega(x,y,t)_{|t\in[10,t_{final}]}$ of the vorticity field upto $t=10$ to $t_{final}=20$. We borrow the same problem setup as in [14], but with less smooth initial conditions (see the initial vorticity in Figure 2), causing higher spatial gradients in the resulting simulation as demonstrated in the first row of Figure 2. We investigate the effectiveness of the proposed framework across different neural operator architectures - Fourier Neural Operator (FNO) [14], UNet, and MATCHO [17]. The FNO accurately learns the dominant low wave-number modes, but remains oblivious to the higher wavenumbers corresponding to the finer spatial features. Additionally, the presence of high spatial gradients in the true solution causes the FNO predictions to exhibit Gibbs oscillations. The UNet and MATCHO also show smoothening in their predictions.
The spectral characteristics of the images generated by diffusion models [44, 45] motivated us to utilize them to improve neural operator predictions. Specifically, we train a score-based diffusion model [35] conditioned on the neural operator’s output, to approximate the ground truth data distribution using denoising score-matching. Once trained, the diffusion model uses annealed Langevin dynamics-based sampling to generate its prediction conditioned on the specific neural operator output. Figure 2 illustrates that the vorticity field estimated by the conditioned diffusion model gets rid of the Gibbs oscillations (for FNO), recovers the finer features, and agrees well with the true power spectrum for all the neural operators compared here. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, FNO, and FNO + Diffusion Model in Figure A.1. The proposed framework demonstrates a better alignment with true modes than the neural operator alone.
Fig.  2: Kolmogorov Flow. The first row shows the vorticity field in a kolmogorov flow simulated using a pseudo-spectral solver. The following rows show the predictions of different types of neural operators - FNO, UNet & MATCHO and that of a score-based diffusion model conditioned on the respective neural operator as a prior. The corresponding energy spectra at each timestep are also displayed.
Fig.  3: Sampling Process. a) An illustration of the sampling process in a score-based diffusion model conditioned on the Fourier neural operator’s prediction (NO + Diffusion), and b) the comparison of the corresponding energy spectra with the ground truth.
2.2 Buoyancy-driven Transport
Here we train a neural operator to learn the buoyancy-driven transport of a scalar concentration field ($d$) [46]. We can define the buoyancy model with Boussinesq approximation using the incompressible Navier-Stokes equation accompanied by the advection transport of $d$ and can be expressed in terms of the buoyancy factor $\xi$ as,
$\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot\nabla\bm{u}$ $=-\frac{1}{\rho}\nabla P+\nu\nabla^{2}\bm{u}+[0,1]^{T}\xi d\quad\text{ s.t }\quad\nabla\cdot\bm{u}=0$ (3)
$\frac{\partial d}{\partial t}+\bm{u}\cdot\nabla d$ $=0$ (4)
where $0\leq t\leq 21$, $\nu=10^{-2}$ and $\xi=0.5$. The system has Dirichlet boundary conditions for the velocity ($\bm{v}=0$) and Neumann boundary conditions for the concentration field ($\frac{\partial d}{\partial x}=0$). We utilized the dataset from PDEArena [47] comprising of 2080 train, 260 validation and 260 test trajectories, generated using $\Phi$flow [48]. Each simulation was saved every $\Delta t=1.5$, resulting in 14 time steps per trajectory.
We train a U-Net as the neural operator to learn the evolution of $d$. The neural operator takes the states at 4 previous timesteps as the input and predicts the next time step as output. During training, the ground truth data is provided as the input. But while validating and testing, the ground truth data at only $t=0,\Delta t,2\Delta t,3\Delta t$ is available. So the neural operator performs an auto-regressive rollout to infer the future states. The setup of the neural operator is consistent with that demonstrated in [47].
In Figure 4, we visualize the ground truth simulation in the first row, predictions of the neural operator in the second row, and that of the diffusion models conditioned on the neural operator’s output in the third row. We compare the spectral distribution of each of the simulations in row 4. We can observe the over-smoothening in the fields predicted by the neural operator. The diffusion model conditioned on the neural operator achieves better spectral similarity with the ground truth, compared to the neural operator alone. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, UNet, and UNet + Diffusion Model in Figure A.2. Again, the proposed framework demonstrates a better alignment with true modes than the neural operator alone.
Fig.  4: Buoyancy-driven Transport. The first row illustrates the true transport of the concentration field $d$ simulated using $\Phi$Flow. The second row is the prediction of the (UNet) neural operator. The third row visualizes the prediction of the diffusion model conditioned on the neural operator’s output. The last row compares the square roots of the respective energy spectra to highlight the differences.
2.3 Turbulent airfoil wake
Here we consider a turbulent wake flow downstream of a NACA0012 airfoil with a chord length of $L_{c}$ at a Reynolds number of 23000, a free stream Mach number of 0.3 and an angle of attack of $6^{\circ}$ [49]. The flow displays coherent structures linked to the Kelvin-Helmholtz instability over the separation bubble and von Karman vortex shedding in the wake, while preserving the broad complexity typical of turbulent flow [50]. This makes it an excellent platform for CFD validation, flow analysis, and the investigation of reduced-complexity models. We directly utilize the large eddy simulation (LES) dataset generated by Towne et al. [50] from "Deep Blue Data", a publicly available database from the University of Michigan. The flow was simulated using a finite-volume based compressible solver - CharLES [51]. The LES dataset is comprised of 16,000 time-resolved snapshots along the mid-span slice, collected at a constant convective time increment of $\tau=0.0104\frac{L_{c}}{U_{\infty}}$. For more details on the problem setup and dataset generation, please refer to section VII in [50].
For training the neural operator we consider a subset of the streamwise velocity component ($u$) of the previously described dataset, consisting of 1000 timesteps, at a time increment of $5\tau$, within a subdomain $4L_{c}\times 2.5L_{c}$. Specifically, we train a U-Net-based neural operator, MATCHO [17], to learn the mapping $u(\bm{x},t)\rightarrow u(\bm{x},t+\Delta t)|_{\Delta t\in[0,50\tau]}$. We split the 1000 timesteps into train, validation and test datasets in the ratio 80:10:10. From Figure 5, we observe that the neural operator predicts a smoothened representation of the true LES simulations. The diffusion model conditioned on the neural operator’s output can recover the finer features, with a better aligned power spectrum. Furthermore, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, MATCHO, and MATCHO + Diffusion Model in Figure A.3. The diffusion model conditioned on the neural operator demonstrates a better alignment with true modes than the neural operator alone.
Fig.  5: Turbulent wake of NACA0012. Distribution of the streamwise velocity $u$ at the mid-span slice at $Re=23,000$. The first row is the ground truth generated from LES, second row is the prediction of neural operator (MATCHO) and the third row is the prediction of a diffusion model conditioned on the neural operator’s output. The last row plots the energy spectrum of the ground truth and predictions.
2.4 3D Turbulent Jet
A jet is a free shear flow consisting of a fast inner stream and a slower outer stream [50]. In a jet, a high-velocity stream of fluid is ejected through a nozzle into the surrounding medium where the mean velocity gradient is not affected by the walls. Simulating the jet facilitates understanding the flow dynamics and optimizing designs in real-world applications in aerospace, automotive and energy industries. Here we try to model a subsonic ($M=0.9$), isothermal, turbulent ($Re\approx 10^{6}$) jet from a convergent-straight nozzle using neural operators.
Again we directly utilize the dataset generated by Towne et al., which is publicly available through the Deep Blue Data repository from the University of Michigan. The simulations were performed with a high-fidelity LES-based compressible flow solver CharLES [51] using the Finite Volume Method. For more details about the problem setup and configurations, please refer section II in [50]. The original dataset consists of 10000 snapshots of the jet sampled every 0.2 acoustic time units $\tau=0.2tc_{\infty}/D$, where $c_{\infty}$ is the free stream speed of sound and $D$ is the nozzle exit diameter.
For training the neural operator we consider the first 1000 timesteps of the velocity field vector ($\bm{u}$) comprising of the axial ($u_{x}$), radial ($u_{r}$) and azimuthal ($u_{\theta}$) components, and utilize the first 800 for training, the next 100 for validation, and the next 100 for testing. The dataset domain is $11.71D$ along the axial direction and $2.78D$ along the radial direction. The neural operator, MATCHO, is trained to learn the mapping $\bm{u}(\bm{x},t)\rightarrow\bm{u}(\bm{x},t+\Delta t)|_{\Delta t\in[0,10\tau]}$.
We illustrate the isosurfaces corresponding to Q-criterion=1.5 in Figure 6. The neural operator is unable to retain the higher wavenumbers in its predictions, especially at the later time states, but captures the general trend of the system contained in the lower wavenumbers. This facilitates the neural operator to behave as an effective prior to the diffusion model which recovers the high frequency components, from angular cross-sections of the velocity vector field inferred by the neural operator. This phenomenon is also illustrated using the power-spectrum plots of the magnitude of the velocity field at different axial cross-sections. Additionally, we perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, MATCHO, and MATCHO + Diffusion Model in Figure A.4. The diffusion model conditioned on the MATCHO demonstrates a better alignment with true modes than the MATCHO alone.
Fig.  6: 3D turbulent jet. Isosurfaces corresponding to Q-criterion = 1.5 of a turbulent jet ($M=0.9$, $Re=10^{6}$) is visualized. The color of the isosurfaces corresponds to the magnitude of the velocity. The first three rows represent solutions of the reference LES, neural operator (MATCHO), and diffusion model conditioned on the neural operator. The last row compares the energy spectra at different axial cross-sections corresponding to $\Delta t=4\tau,8\tau$.
2.5 Schlieren velocimetry of turbulent helium jet in air
So far our results demonstrate the effectiveness of using neural operators as priors of diffusion models, on numerically simulated turbulent databases. In the last test case, we investigate if the proposed framework can be directly applied to real experimental setups. We train the surrogate network to model a turbulent helium jet in air from Schlieren velocimetry data [52].
The system under consideration is a round turbulent helium jet in air expelled from a nozzle of diameter $d$ at a jet exit mach number of $M=0.45$, velocity of $u=436m/s$ and Reynolds number $Re=5890$. The domain of interest is $200d$ along the x-direction and $100d$ along the y-direction. A schematic of the traditional single mirror schlieren apparatus is provided in Figure 7a (screenshot from [52]). A Photron APX-RS camera captures the images of the turbulent jet at 6000 frames/s with an image exposure of 167 $\mu s$.
We directly utilize the publically available schlieren velocimetry dataset provided by [52], in raw .tif format. Each snapshot is separated by $\tau=1/6000s$. For training the neural operator we consider 1000 snapshots of the density gradient field ($f$), and utilize the first 800 for training, next 100 for validation and the last 100 for testing. The neural operator, MATCHO, is trained to learn the mapping $f(\bm{x},t)\rightarrow f(\bm{x},t+\Delta t)|_{\Delta t\in[0,10\tau]}$, and the diffusion model takes the output of MATCHO as its prior.
Figure 7 a) illustrates the experimental setup of a traditional single-mirror schlieren apparatus used in [52]. In part b) the first row represents the density gradient captured using the schlieren velocimetry setup at $\Delta t=4\tau,8\tau$, with nozzle on the left and helium expelled towards right along the x-axis. The second and the third row represents the predictions of neural operator and the diffusion model conditioned on neural operator, again at the same time steps. The color bar represents the density gradient normalized between 0 and 1. In the last row we compare the energy spectra of as a function of wavenumber for the three approaches discussed for both the timesteps. From both the contour plots as well as the energy spectrum, the predictions of diffusion model conditioned on the neural operator align better with the ground truth than the neural operator alone. We also perform POD to analyze the energy decay spectrum and to visualize & compare the POD modes of the Ground Truth, MATCHO, and MATCHO + Diffusion Model in Figure 8. The diffusion model conditioned on the neural operator demonstrates a better alignment with true modes than the neural operator alone.
Fig.  7: Schlieren velocimetry. a) Schematic of knife-based single mirror Schlieren apparatus (screenshot from [52]) Settles et al. uses to visualize turbulent jet of helium in the air, with a jet exit $M=0.45$ and $Re=5890$. b) The density gradients at $\Delta t=4\tau,8\tau$ captured using the schlieren velocimetry apparatus (row 1), estimated by neural operator, MATCHO (row 2) and estimated using diffusion model conditioned on neural operator in (row 3). In the last row, we compare the energy spectra of the respective snapshots captured/estimated by the three approaches.
3 Discussion
Mitigating spectral bias in neural operators
Trained neural operators can serve as fast surrogates for modeling physical systems when subjected to unseen initial or boundary conditions, without the need for solving the system again using traditional discretization-based numerical solvers. However, the states inferred by the neural operator often tend to suffer from over-smoothening. In this study, we systematically compare the distribution of energy across the spectrum of Fourier modes, for the true solution and the states predicted by the neural operator. The energy of the state at any time step can be considered as the square of the absolute of the Fourier transform of the corresponding state. For all the cases reported in section 2, the energy spectrum analysis of the states demonstrates that the energy of the solutions inferred by the neural operator aligns well with the ground truth only for lower wave numbers. For higher wave numbers, the energy of neural operator prediction was consistently lower than that of the true solution. This behavior is called the spectral bias of neural operators and often becomes more prominent when the prediction window horizon is longer. The problem of oversmoothening arising from spectral bias of neural operator becomes a significant drawback while modeling turbulent flows using neural operators, because turbulent systems retains non-trivial energy even at higher wavenumber that cannot be approximated by the neural operators in the conventional setup. Our results demonstrate that a diffusion model conditioned on a neural operator as its prior can overcome this issue and provide estimates of the state with a better aligned energy spectrum.
Computational Costs
The diffusion model has to autoregressively perform $n$ denoising steps to estimate the state of the system at any time step as shown in Figure 3. Recent advances and improvements in the training and sampling of diffusion models [35] helped to bring down the number of sampling steps from $n=1000$ to $n=32$. For the Kolmogorov flow test case, the trained neural operator can predict a trajectory from an unseen initial state in roughly $0.02s$, whereas, the diffusion model requires $0.9s$. Although the diffusion model is slower than the neural operator at the inference, it is significantly faster than the traditional discretization-based solvers and achieves the true spectral properties unlike the neural operators. All the neural operators and the diffusion models consist of $\approx 2M$ trainable parameters. The neural operators were trained using the cosine annealing based learning rate scheduler, and the diffusion models were trained with a constant learning rate of $10^{-4}$ as recommended in [35]. Our diffusion model was implemented based on [35, 53].
Proper Orthogonal Decomposition Analysis
The energy spectra in Figure 2 to Figure 7 demonstrate that the diffusion model conditioned on the neural operator performs better than the neural operator alone. However, these plots are constructed for each time step separately and do not contain information about the cross-correlation of the states across time steps. We address this question by performing proper orthogonal decomposition (POD) on 10 timesteps predicted by the neural operator, as well as the diffusion model conditioned on the neural operator, and compare it with the corresponding state of the ground truth. Figure 8 illustrates the decay of the eigenvalues and compares the POD modes of the true data with the other two approaches for the Schlieren velocimetry case. The proposed framework of the diffusion models conditioned on the neural operator has better aligned eigen spectrum and eigen modes with the ground truth than the neural operators alone. The POD analysis of Kolmogorov flow (Figure A.1), buoyancy-driven flow (Figure A.2), turbulent airfoil (Figure A.3) and 3D turbulent jet (Figure A.4) are also provided in Appendix A.
Fig.  8: POD analysis of Schlieren velocimetry. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator and diffusion model conditioned on neural operator (MATCHO) with the true solution obtained from Schlieren velocimetry. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Future Directions
Summarizing, we propose a framework that utilizes neural operators as effective priors for diffusion models. Our results demonstrate that the proposed framework that blends neural operators with diffusion models significantly improves the energy spectrum compared to the neural operators alone and has been tested across different types of neural operator architectures - FNO, U-Net, and MATCHO, over a variety of test cases. The framework utilized here is general, and therefore, any advances in the quality, computational efficiency, or scalability of either the neural operator or the diffusion models in the future can be applied seamlessly and could potentially improve the proposed framework.
The sampling routine of the score-based diffusion model we adopted consists of 32 denoising steps, corresponding to 32 function evaluations as against 1 function evaluation for the neural operator. Therefore, the diffusion model is approximately 32 times slower than the neural operator at inference. Nevertheless, inferring states of the system using diffusion models still takes only an order of ${}1s$ on an NVIDIA H100 GPU, for all the cases we considered. Improving the accuracy of longer rollouts, incorporating conservation of physical laws, and applying the methodology to problems in other scientific domains are interesting future research directions. We also believe it is worth investigating the potential of combining neural operators with other generative modeling techniques like GANs and normalizing flows.
The utility of this new method is in accelerating DNS and LES by interweaving small bursts of such expensive simulations with large extrapolation time intervals from our method. This is similar to previous work we did for materials problem [19] using the neural operators alone. Moreover, for experimental work where a few snapshots are visualized, the present method will enable a continuous in time evolution of the turbulent field at DNS-level fidelity.
4 Methods
4.1 Neural Operators
Neural Operators are a class of architectures in deep learning, that maps between infinite dimensional functional spaces. Consider an operator $\mathcal{G}$, that maps from the input function $X$ to the output function $Y$, i.e., $\mathcal{G}:X\rightarrow Y$. Then any machine learning architecture that is used to approximate this functional mapping to predict $y\in Y$ for a given $x\in X$ can be defined as a neural operator. Neural operator is a popular framework for encoding dynamics and making predictions under some given conditions. But like any other neural network architectures, they suffer from spectral bias.
Lemma 4.1. Let $F:X\rightarrow Y$ be a functional mapping with $X,Y\in H^{s}(\Omega)$, where $H^{s}(\Omega)$ is the Sobelov space of functions defined on $\Omega$. Let $\mathcal{G}(\theta_{NO})$ be a neural operator parameterized by $\theta_{NO}$, that is approximating the functional space $F$. Then $\mathcal{G}$ learns the lower frequency components of the functional space $F$ more effectively.
Sketch of Proof: We want to approximate the functional space $F(X)=Y$, then if we take the Fourier transform for this function, we have
[…]
and
[…]
where $\hat{F}(\xi)$ represents the amplitude of the the frequency component $\xi$. Now let $\mathcal{G}(\theta_{NO},X)$ is the operator approximating $F(X)$, typically the loss function $\mathcal{L}(\theta_{NO})$ used for optimizing such networks is given by
$\mathcal{L}(\theta_{NO})=\frac{1}{N}\sum_{n=1}^{N}(F(x_{n})-\mathcal{G}(\theta_{NO};x_{n}))^{2},$ (7)
where $x_{n}\in X$, and $n=1,2,\cdot\cdot\cdot,N$ are the discrete points where the functional evaluation are available for training. We can also express $\mathcal{G}$ in terms of its discrete Fourier components as
$\hat{\mathcal{G}}(\theta_{NO};\xi)=\frac{1}{N}\sum_{n=1}^{N-1}\mathcal{G}(\theta_{NO};x_{n})e^{-2\pi i\xi n/N},$ (8)
and
$\mathcal{G}(\theta_{NO};x_{n})=\frac{1}{N}\sum_{n=1}^{N-1}\hat{\mathcal{G}}(\theta_{NO};\xi)e^{2\pi i\xi n/N}.$ (9)
During the minimization of the loss functions the parameters are updated based on the gradient descent update rule, given by
$\theta_{NO}^{(m+1)}=\theta_{NO}^{(m)}-\alpha\nabla_{\theta_{NO}}\mathcal{L}(\theta_{NO}^{(m)}),\hskip 14.45377ptm=1,2,\cdot\cdot\cdot,M$ (10)
where $M$ is the total number of iterations, $\alpha$ is the learning rate and $\nabla_{\theta_{NO}}\mathcal{L}$ is given by
$\nabla_{\theta_{NO}}\mathcal{L}(\theta_{NO})=\frac{-2}{N}\sum_{n=1}^{N}(F(x_{n})-\mathcal{G}(\theta_{NO};x_{n}))\nabla_{\theta_{NO}}\mathcal{G}(\theta_{NO};x_{n}).$ (11)
Hence, mathematically the rate of change of the Fourier coefficients $\hat{\mathcal{G}}(\theta_{NO};\xi)$ with respect to training iteration $m$ can be given by
$\frac{d}{dm}(\hat{\mathcal{G}}(\theta_{NO};\xi))\propto-\alpha\hat{\nabla}_{\theta_{NO}}\mathcal{L}(\theta_{NO}).$ (12)
$\alpha\hat{\nabla}_{\theta_{NO}}\mathcal{L}(\theta_{NO})$ is larger for smaller $\xi$. Therefore, the operator $\mathcal{G}$ learns the low frequency components of the functional space $F(X)$ more effectively compared to higher frequency components, which is known as the spectral bias. Hence like any other networks, operators suffer from spectral bias, which leads to ineffective refinement of the higher frequency component, leading, in turn, to sharp decay of the power spectrum associated with the output predicted by any neural operator. Although neural operators fail to capture the high frequency, it captures effectively the mean flow dominated by the low wave numbers. This acts as an ideal prior for diffusion models to mimic the true flow consisting of all the frequencies.
4.2 Score-based diffusion models
The overall goal of a score ($s$) based diffusion model is to estimate the score function defined by $s_{\theta_{D}}(\mathbf{X})=\nabla_{X}\log p(\mathbf{X})$, where $\theta_{D}$ are the parameters of the diffusion model and $p$ is the probability density of $\mathbf{X}$. Since the data distribution is not known explicitly, and there is a chance that the data might lie on a lower dimensional manifold, the score function may be ill defined in regions where there is no data. To circumvent this issue, perturbation of the data is done by adding Gaussian noise, which ensures that the score is well defined across the entire space by smoothing the distribution. Hence, the score function gives us a direction to move towards the region of higher probability. However the direct way to sample from the distribution is missing here. In order to fix this problem, Langevin Dynamics was used in [54]. Langevin Dynamics ensures the convergence of the generated samples to the true underlying distribution. It balances the inherent dynamics between deterministic movement, which is driven by the gradient of log of the probability, with random exploration given by the noise levels. In our proposed method, we use the score function, which is additionally conditioned on the output of the neural operator $\mathbf{Y}=\mathbf{\mathcal{G}}(\theta_{NO},\mathbf{X})$ given by
$s_{\theta_{D}}(\mathbf{X},\sigma,\mathbf{Y})=\nabla_{X}\log p(\mathbf{X}|\mathbf{Y},\sigma),$ (13)
where $\sigma$ is the noise level. The modified score function now considers the joint distribution of $\mathbf{X}$ and $\mathbf{Y}$. This ensures that the the generated samples from the diffusion model are consistent with both the structures defined by $\mathbf{Y}$ and the data distribution that is learned by the diffusion model. The update rule for Langevin dynamics can be shown as
$\mathbf{X}_{i+1}=\mathbf{X}_{i}+\frac{\varepsilon}{2}s_{\theta_{D}}(\mathbf{X}_{i},\sigma_{t},\mathbf{Y})+\sqrt{\varepsilon}z_{i},$ (14)
where $\varepsilon$ is the step size and $z_{i}$ is the noise component. From Eqns 13 and 14, we can see that as $\sigma_{i}$ decreases, the score function $s_{\theta_{D}}$ focuses on recovering the high-frequency details which was smoothen out by the noise. In the frequency domain Eqn 14 can be shown as
$\hat{\mathbf{X}}_{i+1}=\hat{\mathbf{X}}_{i}+\frac{\varepsilon}{2}\hat{s}_{\theta_{D}}(\mathbf{X}_{i},\sigma_{i},\mathbf{Y})+\sqrt{\varepsilon}\hat{z}_{i},$ (15)
where $\hat{\mathbf{X}}_{i}$, $\hat{s}$ and $\hat{z}$ are the Fourier transform of the respective functions. Hence, after multiple iterations as the noise diminishes $\hat{z}$ becomes negligible and the score function amplifies the high frequency component as it acts as a high pass filter in reverse.
Lemma 4.2. Let $u(\mathbf{X},t)$in $H^{s}(\Omega):\{u\in L^{2}(\Omega)|\partial^{\alpha}u\in L^{2}(\Omega),\forall|\alpha|\leq s\}$ be the true solution discretized over N spatial grid points.
$u(\mathbf{X},t)=\sum_{n=1}^{\infty}a_{n}(t)\phi_{n}(\mathbf{X}),$ (16)
where $\phi_{n}(\mathbf{X})$ are the basis functions and $a_{n}(t)$ are the time-dependent coefficients. Let $U_{NO}(\mathbf{X},t)$ and $U_{D}(\mathbf{X},t)$ represents the output of the neural operator and the diffusion model respectively, such that
$U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t)\simeq\sum_{n=1}^{\infty}a_{n}(t)\phi_{n}(\mathbf{X}).$ (17)
If $k_{\xi}$ represents the wave number which determines the basis function approximated by $U_{NO}$ such that
$E_{U_{NO}}(k_{i})=\frac{1}{N}\left|\sum_{(k_{n}\leq k_{\xi})}a_{n}(t)\hat{\phi_{n}}(k_{i})\right|^{2},$ (18)
where $E_{U_{NO}}(k_{i})$ is the energy spectrum related to $U_{NO}$ in the discrete Fourier domain at wave number $k_{i}$, then $\left|u(\mathbf{X},t)-(U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t))\right|\rightarrow 0$ as $k_{\xi}$ increases.
Sketch of Proof: We have,
$U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t)$ $\simeq\sum_{n=1}^{\infty}a_{n}(t)\phi_{n}(\mathbf{X})$ (19)
$\implies U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t)$ $=\sum_{n:k_{n}\leq k_{\xi}}a_{n}(t)\phi_{n}(\mathbf{X})+\sum_{n:k_{n}>k_{\xi}}a_{n}(t)\phi_{n}(\mathbf{X}).$ (20)
In the discrete Fourier space we can write it in terms of the energy spectrum as
$E_{U_{NO}}(k_{i})+E_{U_{D}}(k_{i})=\frac{1}{N}\left|\sum_{(k_{n}\leq k_{\xi})}a_{n}(t)\hat{\phi_{n}}(k_{i})\right|^{2}+\frac{1}{N}\left|\sum_{(k_{n}>k_{\xi})}a_{n}(t)\hat{\phi_{n}}(k_{i})\right|^{2}.$ (21)
From Eqn 13 in the frequency domain, the score function can be written as with the assumption that the conditional distribution $p$ follows a Gaussian distribution
$\hat{s}(k_{i},U_{NO},\sigma_{j})\simeq\nabla_{\hat{u}(k_{i})}\log p(\hat{u}(k_{i})|\hat{U}_{NO}(k_{i}),\sigma_{j})=-\frac{\hat{u}(k_{i})-\hat{U}_{NO}(k_{i})}{\sigma^{2}(k_{i},\sigma_{j})},$ (22)
where $\hat{u}(k_{i})$ is the Fourier transform of the current state at wave number $k_{i}$ and $\sigma^{2}(k_{i},\sigma_{j})$ is the variance of the distribution conditioned on the noise level $\sigma_{j}$ at wave number $k_{i}$. Although the specific relation between $\sigma^{2}(k_{i},\sigma_{j})$ and $k_{i}$ is problem specific, typically [55],
$\sigma^{2}(k_{i},\sigma_{j})\propto\frac{1}{k_{i}^{p}}+\sigma_{j}^{2},$ (23)
where $p$ is a positive constant based on the data. Hence, as $k_{i}\rightarrow 0$, $\sigma^{2}(k_{i},\sigma_{j})$ increases and $\frac{\hat{u}(k_{i})-\hat{U}_{NO}(k_{i})}{\sigma^{2}(k_{i},\sigma_{j})}\rightarrow 0$. This implies a weak gradient value signifying smaller steps. Now if we look at the frequency domain of the Langevin dynamics, from Eqn 15
$\hat{u}_{j+1}=\hat{u}_{j}(k_{i})-\frac{\varepsilon}{2}\frac{\hat{u}(k_{i})-\hat{U}_{NO}(k_{i})}{\sigma^{2}(k_{i},\sigma_{j})}+\sqrt{\varepsilon}\hat{z}_{j}(k_{i}),$ (24)
it can be seen that for smaller $k_{i}$ the above equation changes to
$\hat{u}_{j+1}\simeq\hat{u_{j}}(k_{i})+\sqrt{\varepsilon}\hat{z}_{j}(k_{i}).$ (25)
This signifies that the Langevin dynamics updates are dominated by the noise term $\sqrt{\varepsilon}\hat{z}_{j}(k_{i})$ instead of the gradient driven term and hence leads to ineffective refinement of the low wave numbers. Hence, in an intermediate step during the diffusion process when the updates are dominated by the noise term, the energy spectrum can be given as
$E_{U_{D,j+1}}(k_{i})=\frac{|\hat{u}_{j}(k_{i})|^{2}}{N}+\frac{\varepsilon|\hat{z}_{j}(k_{i})|^{2}}{N}+\frac{2\sqrt{\varepsilon}\Xi[\hat{u}_{j}(k_{i})\hat{z}^{\ast}_{j}(k_{i})]}{N},$ (26)
where $\Xi[.]$ represents the real part of the complex number and $\hat{z}^{\ast}$ the complex conjugate of $\hat{z}$. Due to weaker gradients, the term $|\hat{u}_{j}(k_{i})|^{2}$ grows slowly and $\varepsilon|\hat{z}_{j}(k_{i})|^{2}$ may dominate, which might lead to flatter energy spectrum at low wave number and may not accurately represent the large scale features. Therefore, as as $k_{\xi}$ increases $\left|u(\mathbf{X},t)-(U_{NO}(\mathbf{X},t)+U_{D}(\mathbf{X},t))\right|\rightarrow 0$, or in other words for the proposed method to work effectively, the $U_{NO}(\mathbf{X},t)$ should be at-least be able to capture the mean flow, in order for the diffusion model to be effectively helping in adding the high frequency features to the predicted mean flow.
In summary, based on Lemma 4.1. and Lemma 4.2. we propose our method where we use a neural operator as a prior to score-based diffusion model. In step 1 we train a neural operator that accurately captures the low frequency wave numbers and hence the mean flow dynamics. In step 2 we pass the neural operator predicted outputs to the diffusion model, which acts a prior, which we further train based on the score function that focuses on the joint distribution and the high frequency components are added which leads to the recovery of the full spectrum. We note that as shown in Lemma 4.2., the neural operator should be able to capture the mean flow accurately for the proposed method to work effectively.
Acknowledgements
Funding: This work was supported by the MURI grant FA9550-20-1-0358 and the the ONR Vannevar Bush Faculty Fellowship (N0001422-1-2795).
Author Contributions:
V.O.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing and Visualization.
A.B.: Conceptualization, Methodology, Software, Formal Analysis, Investigation, Writing - Original Draft and Writing - Review & Editing
Z.Z.: Software, Validation, Formal Analysis, Investigation and Visualization
G.E.K.: Validation, Formal Analysis, Resources, Writing - Review & Editing, Supervision, Project administration and Funding acquisition
Competing Interests: The authors declare that they have no competing interests.
Data and Materials Availability: The data for Kolmogorov flow was simulated using the publicly available pseudo-spectral solver provided in FNO’s repository. The buoyancy-driven transport data was from PDEArena. The LES datasets for turbulent airfoil and jet were downloaded from link. The Schlieren velocimetry dataset can be found here. The code developed in this study will be made available after publication in our GitHub repository.
Appendix A POD Analysis
In the main manuscript, we investigate the energy spectra with respect to the wavenumber corresponding to the neural operator and diffusion model conditioned on the neural operator, computed on each time step separately. Here we perform a detailed proper orthogonal decomposition (POD) analysis by performing singular value decomposition on 10 consecutive timesteps of the trajectory, and plot the eigen energy spectrum. Additionally, we also visualize the corresponding eigen modes/basis required for accurate reconstruction of the aforementioned timesteps. The POD analysis for Kolmogorov flow Figure A.1, buoyancy driven transport Figure A.2, turbulent airfoil Figure A.3, 3D turbulent jet Figure A.4 and schlieren velocimetry Figure 8 are shown below.
Fig.  A.1: POD analysis of Kolmogorov flow. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (FNO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.2: POD analysis of buoyancy-driven flow. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (UNet) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.3: POD analysis of a turbulent airfoil. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (MATCHO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.
Fig.  A.4: POD analysis of a 3D jet We perform POD on the angular velocity component at $\theta=0$. We compare the decay of eigenvalues obtained from performing proper orthogonal decomposition of 10 snapshots predicted by the neural operator (MATCHO) and diffusion model conditioned on the neural operator with the ground truth solution. Additionally, we also visualize and compare POD modes corresponding to $i=1,3,5,7,9$.